At t = 0 s, a particle starts from the origin of a coordinate system and moves in the xy plane with a velocity ~v = (7.9ˆi − 3.2ˆj) m/s. Determine the x position of the particle at t = 2.0 s.
Answer #1

V = dX/dt

=> dX = Vdt = (7.9ˆi − 3.2ˆj)*dt

integrating both side

we get

Xfinal - Xinitial = (7.9tˆi − 3.2tˆj)

Xintial = 0 , at t=0

=> Xfinal at t =2

Xfinal = 7.9*2 i - 3.2*2j = 15.8i - 6.4j